At Ed Helper coffee, one kind of coffee sells for $1 a cup and another sells for $3.25 a cup. How much of each coffee should be used to make 9 cups of a coffee mixture which sells for $2.75 a cup?
Answer by mananth(16949)

--------Price ------------ Numbers
Coffee I 1 ----------------x
Coffee II 3.25 ----------------9-x
Total 2.75 ----------------9
x+3.25(9-x)=2.75*9
x+29.25-3.25x=24.75
x-3.25x =24.75-29.25
-2.25x= -4.5
/-2.25
x=2 ===>$1.00
7====>3.25%